OpenStudy (anonymous):

The cost of a high school ring is $145 for the large size and $115 for the regular size. The total receipts form the sale of 120 rings was $15,060. How many rings of each size were sold?

OpenStudy (ganpat):

let x be large size and y = regular size.. so, you have two equation.. 145x + 115y = 15,060 and x + y = 120.. can u solve them ??

OpenStudy (anonymous):

145+115=260 so it would be: 260=15,060 I don't think i'm doing this right.

OpenStudy (ganpat):

u r going wrong way.. x + y = 120 x = 120 - y 145x + 115y = 15,060 140 (120 - y) +115y = 15060 16800 - 140y + 115y = 15060 25y = 1740 y = 69.6 = 70 x = 120 - 70 = 50...

OpenStudy (anonymous):

So 50 large rings were sold and 70 small were sold?

OpenStudy (anonymous):

right?